OpenStudy (anonymous):

solve the equation (x^2-23/2x^2-5x-3)+(2/x-3)= (1/2x+1)

OpenStudy (anonymous):

should do some more brackets, cant tell whats where, from that input i say 4 solutions, 2 complex

OpenStudy (dumbcow):

no it factors \[\rightarrow \frac{(x+7)(x-3)}{(2x+1)(x-3)} = \frac{1}{2x+1}\] \[\rightarrow x+7 = 1\] \[x = -6\]

OpenStudy (anonymous):

so much more proper when i can read it how it should be written. nice